OpenStudy (anonymous):

Solve S = 4v2 for v. The answer I got was 2/s = v. I have no idea what I did wrong. Please help?

OpenStudy (australopithecus):

S = 4v^2 S/4 = v^2 (S/4)^1 = v^(2) \[(S/4)^{(1)(\frac{1}{2})} = v^{(2)(\frac{1}{2})}\] \[(S/4)^{(\frac{1}{2})} = v^{1}\] \[(S/4)^{\frac{1}{2}} = v\] or \[\sqrt{\frac{S}{4}} = v\]

OpenStudy (australopithecus):

or \[\frac{\sqrt{S}}{2} = v\]

OpenStudy (anonymous):

Okay. Thanks tons. :)

OpenStudy (australopithecus):

wait one sec \[\pm \frac{\sqrt{S}}{2} = v\]

OpenStudy (anonymous):

Could you explain the steps a bit more...?

OpenStudy (australopithecus):

that would be the actual answer

jimthompson5910 (jim_thompson5910):

If v is strictly positive or zero, then \[\Large v = \frac{\sqrt{S}}{2}\] But if there is no restriction, then \[\Large v = \pm\frac{\sqrt{S}}{2}\]

OpenStudy (australopithecus):

it is plus or minus because you are dealing with a square function thus you can input a negative of postive value into it and still get a positive

OpenStudy (australopithecus):

what dont you understand about the algebra?

OpenStudy (australopithecus):

I dont think I could write it in anymore detail

OpenStudy (australopithecus):

The only thing I think you could get confused about is the exponents

jimthompson5910 (jim_thompson5910):

Basic idea: Step 1) Divide both sides by 4 Step 2) Take the square root of both sides Step 3) Simplify

OpenStudy (australopithecus):

jim thompson that might be the basic idea but it doesn't get to the root of the problem learning how to deal with exponents is fundamental

jimthompson5910 (jim_thompson5910):

no need to deal with exponents really

OpenStudy (australopithecus):

if he had v^(3) or v^(3/2) he wouldnt be able to solve for it

jimthompson5910 (jim_thompson5910):

sure you can go that route, but dealing with square roots is probably easier

OpenStudy (australopithecus):

Why do you bother even posting if you aren't going to educate people properly

jimthompson5910 (jim_thompson5910):

there's nothing wrong with my solution

OpenStudy (australopithecus):

no there isn't but this site isn't about solving peoples problems its about teaching them how to solve their own problems

jimthompson5910 (jim_thompson5910):

I do that, i post steps so they can follow them and learn from them

OpenStudy (australopithecus):

yeah I dont know why Im arguing with you sorry person

jimthompson5910 (jim_thompson5910):

and i go over how I got those steps and such

jimthompson5910 (jim_thompson5910):

yeah idk either, you're very rude

OpenStudy (australopithecus):

yeah and I'm a lot of other stuff too :)

jimthompson5910 (jim_thompson5910):

I guess this is the thanks I get for helping people...

OpenStudy (australopithecus):

do you want me to mail you a sandwich

jimthompson5910 (jim_thompson5910):

i don't need your sarcasm, and if you haven't noticed I provide a lot of hints and such where I don't do the full problem

jimthompson5910 (jim_thompson5910):

or I'll do a piece of it and ask them to finish or walk me through it

OpenStudy (australopithecus):

It isn't sarcasm give me your address and I will mail you a sandwich

OpenStudy (australopithecus):

I'm not judging you man, sorry I said anything stop being overly sensitive

jimthompson5910 (jim_thompson5910):

well i would appreciate you not judging or jumping to conclusions, I'm not being overly sensitive

OpenStudy (australopithecus):

I would appreciate you not jumping to conclusions about me and sarcasm but life isn't the way we always want it to be is it

jimthompson5910 (jim_thompson5910):

no, it isn't...we offer our help only to have people snap back, but oh well

OpenStudy (australopithecus):

meh